OpenStudy (anonymous):

Rene is constructing two similar boxes shaped as similar rectangular prisms. The smaller box has a height of 12 inches and a square base of side length 8 inches. If the larger box has a height of 15 inches, what is its volume? Answer 960 cubic inches 1200 cubic inches 1440 cubic inches 1500 cubic inches

OpenStudy (campbell_st):

let x be the base length of the larger box using rations for similar shapes 15/12 = x/8 x = (15 x 8)/12 calculate the side length then find the base area and multiply by the height of 15 to find the volume

OpenStudy (anonymous):

so the answer would be?

OpenStudy (anonymous):

1500?

OpenStudy (campbell_st):

looks like it to me

OpenStudy (anonymous):

thank you